# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

### v1

@@ 1.2 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
Under assumption ($A_{0}$), the spectrum of each $\mathscr{A}_{m,j}$ consists of the nonzero eigenvalue
$\mu_{m,j}:=1+\frac{\beta^{2}\mu_{j}-1}{1+m^{2}}=\frac{m^{2}+\beta^{2}\mu_{j}}{1+m^{2}},$
with multiplicity $m_{j}\in{\mathbb{N}}$ (cf. discussion of isotypic multiplicity (6)). The product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix A) permits us to express the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ in terms of a Burnside Ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{m,j}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j}$ of the $G$-equivariant linear isomorphism $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-isotypic components $\mathscr{H}_{m,j}$ on their respective open unit balls $B(\mathscr{H}_{m,j}):=\{u\in\mathscr{H}_{m,j}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@
1.3 Orbit Types of Maximal Kind
@@
One of the principal difficulties when working with a product group of the form $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ is that the conjugacy class of the subgroup $O(2)\times\{e_{\Gamma}\}\times\{1\}\leq G$ is always a generator of the Burnside Ring $A(G)$. Since any orbit type $(H)$ which is maximal in the isotropy lattice $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must also be maximal in sublattice $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$, this implies that any non-zero function $u\in\mathscr{H}$ with an isotropy group $G_{u}\leq G$ such that $(G_{u})$ is a maximal element in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be $O(2)$-invariant. Therefore, any map $u\in\mathscr{H}$ which is both non-trivial and non-constant has an isotropy group $G_{u}\leq G$ with a conjugacy class that is maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m$.
@@
At this point, we can strengthen the sufficient condition for the existence of a non-trivial solution to (2) presented in Section 1.1 to a sufficient condition for the existence of a non-stationary solution; namely, if for some orbit type of maximal kind $(H)\in\mathfrak{M}$ one has
@@ 1.4 The $s$-Folding Homomorphism
1.5 An Outline of the Paper
The remainder of this paper is organized as follows: In Section 2, we present a classification for the generators of the Burnside Ring $A(G)$ according to their lattice relations with respect to the $s$-folding homomorphism (12). This classification is then used, in Section 3, to derive a practical formula for characterizing the non-triviality of an orbit type of maximal kind in the product of a finite number of basic degrees. Specifically, we find that the contribution of each eigenvalue $\mu_{m,j}\in\sigma(\mathscr{A})$ to the appearance of an orbit type of maximal kind $(H)\in\mathfrak{M}$ in the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ depends on $(i)$ the sign of $\mu_{m,j}$ and $(ii)$ the parity of the dimension of the corresponding $H$-fixed-point space $\mathcal{V}_{m,j}^{H}$. To formalize these findings, we introduce the following definition:
Definition 1.4.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. An eigenvalue $\mu_{j}\in\sigma(A)$ is called $H$-essential if the following two relations are satisfied
$\mu_{j}<\frac{-m^{2}}{\beta^{2}},\quad 2\nmid\dim\mathcal{V}_{m,j}^{H}.$
Remark 1.5.
In the field of equivariant bifurcation theory, the Ize Conjecture (IC) refers to the proposition that every irreducible representation of a compact Lie group admits at least one subgroup with an odd dimensional fixed point space. Although the IC was shown to be false by Lauterbach et al. in [12] with counter examples of irreducible representations which only admit even dimensional fixed point spaces, the conjecture holds in many cases.
Finally, in Section 4, we demonstrate how these results lead to our main existence theorem for the system (1):
Theorem 1.6.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. If the matrix $A:V\rightarrow V$ admits an odd number of $H$-essential eigenvalues, then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
@@ 1.5 An Outline of the Paper
Since, the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside Ring products of the form (10). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the Leray-Schauder $G$-equivariant degree (cf. Appendix A) implies

### v2

@@ 2.1 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
Under assumption ($A_{0}$), the spectrum of each matrix $\mathscr{A}_{m,j}$ consists of the nonzero eigenvalue
$\mu_{m,j}:=1+\frac{\beta^{2}\mu_{j}-1}{1+m^{2}}=\frac{m^{2}+\beta^{2}\mu_{j}}{1+m^{2}},$ (7)
with multiplicity $m_{j}\in{\mathbb{N}}$ (cf. discussion of isotypic multiplicity (2)). The product property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A) permits us to express the degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$ in terms of a Burnside Ring product of the $G$-equivariant Leray-Schauder degrees of the various restrictions $\mathscr{A}_{m,j}:\mathscr{H}_{m,j}\rightarrow\mathscr{H}_{m,j}$ of the $G$-equivariant linear isomorphism $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-isotypic components $\mathscr{H}_{m,j}$ on their respective open unit balls $B(\mathscr{H}_{m,j}):=\{u\in\mathscr{H}_{m,j}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@
3 Orbit Types of Maximal Kind
@@
One of the principal difficulties when working with a product group of the form $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ is that the conjugacy class of the subgroup $O(2)\times\{e_{\Gamma}\}\times\{1\}\leq G$ is always a generator of the Burnside Ring $A(G)$. Since any orbit type $(H)$ which is maximal in the isotropy lattice $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must also be maximal in the sublattice $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$, this implies that any non-zero function $u\in\mathscr{H}$ with an isotropy group $G_{u}\leq G$ such that $(G_{u})$ is a maximal element in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be $O(2)$-invariant. Therefore, any map $u\in\mathscr{H}$ which is both non-trivial and non-constant has an isotropy group $G_{u}\leq G$ with a conjugacy class that is maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m$.
@@
At this point, we can strengthen the sufficient condition for the existence of a non-trivial solution to (4) stated in Lemma 1.1 to a sufficient condition for the existence of a non-stationary solution:
Lemma 3.2.
If for some orbit type of maximal kind $(H)\in\mathfrak{M}$ one has
@@ 3.1 The $s$-Folding Homomorphism
Since, the classification put forward in the previous section fully characterizes the orbit types $(\tilde{H})$ with $(\tilde{H})>(H)$ for any orbit type of maximal kind $(H)\in\mathfrak{M}$, we are now in a position to calculate their coefficients in Burnside Ring products of the form (11). Given a basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ corresponding to an index pair $(m,j)\in\Sigma$ with $m>0$ and an orbit type of maximal kind $(H)\in\mathfrak{M}$, the recurrence formula for the $G$-equivariant Leray-Schauder degree (cf. Appendix A) implies
@@ 6 On the Non-Triviality of the Degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
7 Motivating Example: An Arrangement of $N$ Coupled Pendula with Dihedral Symmetries and Subject to Nonlinear Forcing
Figure 1: Cycle of $N$ oscillating pendula with $\Gamma=D_{N}$ symmetries
We can provide a more definitive prediction of the symmetries expressed by non-stationary solutions to (1) than that established in Theorem 1.2 by prescribing an exact coupling relation between the pendula in our configuration. Specifically, let $\Gamma$ be the dihedral group of order $2N$ such that our full symmetry group becomes
$G:=O(2)\times D_{N}\times{\mathbb{Z}}_{2}.$
Our model for this configuration is the system
$\begin{cases}\ddot{x}(t)=|u|^{q}u-(L+\operatorname{Id})u,\quad t\in\mathbb{R},\;x(t)\in V\\ x(t)=x(t+p),\;\dot{x}(t)=\dot{x}(t+p),\end{cases}$ (32)
where $|u|^{q}u:=(|u_{1}|^{q}u,|u_{2}|^{q}u,\ldots,|u_{N}|^{q}u)$ for any even $q>1$ and $L:V\rightarrow V$ is the graph Laplacian matrix for an undirected graph $\mathbb{G}$ invariant under the permutation action of $D_{N}$, with $N$ vertices representing a collection of the same number of oscillating pendula and with edges representing the coupling relations between vertices. Since the natural permutation representation of $\Gamma$ on $V$
$\rho_{V}:\Gamma\rightarrow GL(V),\quad\rho_{V}(\sigma)(u_{1},u_{2},\ldots,u_{N}):=(u_{\sigma(1)},u_{\sigma(2)},\ldots,u_{\sigma(N)}),$ (33)
is induced by the permutation actions of the rotation and reflection generators $\gamma,\kappa\in\Gamma$ on the indices $i\in\{1,\ldots,N\}$
$\gamma(i):=i+1\pmod{N}\text{ and }\kappa(i):=N-i\pmod{N},$
its character can be determined according to the rule
$\chi_{V}(\gamma)=0\;\text{ and }\;\chi_{V}(\kappa)=\begin{cases}1&\text{ if }N\text{ is odd};\\ 0&\text{ if }N\text{ is even}.\end{cases}$
The number and character of the irreducible $\Gamma$ representations also depend on the dihedral order $N$: there are always the trivial representation $\mathcal{V}_{0}\simeq\mathbb{R}$, on which $\Gamma$ acts trivially, $\lfloor\frac{N+1}{2}\rfloor-1$ geometric representations $\mathcal{V}_{j}$, each with an action induced by the corresponding matrix representation $\rho_{j}:\Gamma\rightarrow GL({\mathbb{C}})$
$\rho_{j}(\kappa):=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\quad\rho_{j}(\gamma):=\begin{pmatrix}\cos(\frac{2j\pi}{N})&-\sin(\frac{2j\pi}{N})\\ \sin(\frac{2j\pi}{N})&\cos(\frac{2j\pi}{N})\end{pmatrix},\quad 1\leq j<\lfloor\frac{N+1}{2}\rfloor,$
the sign representation $\mathcal{V}_{*}\simeq\mathbb{R}$, with the action
$\sigma\cdot v:=\text{\rm sign\,}(\sigma)v,\quad v\in\mathcal{V}_{*},$
and, in the case that $N$ is even, two additional irreducible one-dimensional representations $\mathcal{V}_{\lfloor\frac{N+1}{2}\rfloor},\mathcal{V}_{**}\simeq\mathbb{R}$ equipped, respectively, with the actions
$\sigma\cdot v:=-\text{\rm sign\,}(\sigma)v,\quad v\in\mathcal{V}_{\lfloor\frac{N+1}{2}\rfloor},$
and
$\rho_{**}(\kappa)=\rho_{**}(\gamma)=-1.$
Comparing characters for the irreducible representations of $\Gamma$ with the character of $V$
conjugacy classes $e_{\Gamma}$ $\kappa$ $\gamma$
$\chi_{0}$ $1$ $1$ $1$
$\chi_{1}$ $2$ 00 $\cos(\frac{2\pi}{N})+\cos(\frac{2\pi}{N})$
$\vdots$ $\vdots$ $\vdots$ $\vdots$
$\chi_{j}$ $2$ 00 $\cos(\frac{2j\pi}{N})+\cos(\frac{2j\pi}{N})$
$\vdots$ $\vdots$ $\vdots$ $\vdots$
$\chi_{\lfloor\frac{N+1}{2}\rfloor}$ $1$ $-1$ $1$
$\chi_{*}$ $1$ $1$ $-1$
$\chi_{**}$ $1$ $-1$ $-1$
$\chi_{V}$ $N$ $\frac{1}{2}(1-(-1)^{N})$ 00
Table 1: Character Table for $D_{N}$.
one obtains the relation
$\chi_{V}=\begin{cases}\chi_{0}+\chi_{1}+\cdots+\chi_{\frac{N+1}{2}-1}&\text{if }2\nmid N;\\ \chi_{0}+\chi_{1}+\cdots+\chi_{\frac{N}{2}-1}+\chi_{\frac{N}{2}}&\text{if }2\mid N,\end{cases}$ (34)
implying that $V$ has the $\Gamma$-isotypic decomposition
$V=\begin{cases}V_{0}\oplus V_{1}\oplus\cdots\oplus V_{\frac{N+1}{2}-1}&\text{if }2\nmid N;\\ V_{0}\oplus V_{1}\oplus\cdots\oplus V_{\frac{N}{2}-1}\oplus V_{\frac{N}{2}}&\text{if }2\mid N.\end{cases}$ (35)
For the sake of generality, we adopt the following notation for the set of $\Gamma$-isotypic indices relevant to the $\Gamma$-isotypic decomposition of $V$
$\mathfrak{J}(N):=\begin{cases}\{0,1,\ldots,\frac{N+1}{2}-1\}&\text{if }2\nmid N;\\ \{0,1,\ldots,\frac{N}{2}-1,\frac{N}{2}\}&\text{if }2\mid N.\end{cases}$
The graph Laplacian associated with the undirected graph $\mathbb{G}$ with $N$ vertices invariant under the permutation action of $D_{N}$ has the form
$L=\left(\begin{array}[c]{ccccccc}-2&1&0&\dots&0&0&1\\ 1&-2&1&\dots&0&0&0\\ 0&1&-2&\dots&0&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&0&\dots&-2&1&0\\ 0&0&0&\dots&1&-2&1\\ 1&0&0&\dots&0&1&-2\end{array}\right).$
One can verify that such a matrix has the eigenvectors
$v_{j}:=(1,\gamma^{j},\gamma^{2j},\ldots,\gamma^{(N-1)j})^{T},\quad j\in\{0,1,\ldots,\lfloor\frac{N+1}{2}\rfloor-1\},\;\gamma:=e^{\frac{2\pi i}{N}}$
corresponding to the eigenvalues
$-z_{j}:=-2+\gamma^{j}+\gamma^{-j}=-4\sin^{2}\left(\frac{\pi j}{N}\right).$
And, in the case that $N$ is even, one will find that there is also the eigenpair
$v_{\frac{N}{2}}:=(1,-1,1\ldots,-1)^{T},\quad z_{\frac{N}{2}}=-4.$
Since each isotypic component of $V$ has simple isotypic multiplicity, i.e. since
$m_{j}=1,\quad j\in\mathfrak{J}(N),$
the eigenvalues of $\mathscr{A}$ (cf. Definition (7)) have the form
$\mu_{m,j}=\frac{m^{2}-\beta^{2}(z_{j}+1)}{1+m^{2}}.$
Therefore, the condition $\mu_{m,j}<0$ is equivalent to the condition
$14\sin^{2}\left(\frac{\pi j}{N}\right)>\frac{m^{2}}{\beta^{2}}+$ 1
and Theorem 1.2 becomes:
Theorem 7.1.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. For the symmetry group $\Gamma=D_{N}$ and under the assumptions ($A_{0}$)($A_{4}$), if the relations
$4\sin^{2}\left(\frac{\pi j}{N}\right)>\frac{m^{2}}{\beta^{2}}+1\text{ and }2\nmid\dim\mathcal{V}_{m,j}^{H},$
are satisfied for an odd number of $D_{N}$ isotypic indices $j\in\mathfrak{J}(N)$, then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (32) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
7.1 The Special Case of $\Gamma=D_{8}$
While the framework described in Section 7 applies to any number of pendula, we can explicitly identify spatio-temporal symmetries of the non-stationary solutions to (32) by choosing any particular dihedral order. Let’s examine the special case of $8$ pendula, such that our full symmetry group becomes
$G:=O(2)\times D_{8}\times{\mathbb{Z}}_{2}.$
Since the non-triviality of the coefficient standing next to an orbit type of maximal kind $(H)\in\mathfrak{M}_{m}$ in the basic degree $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is equivalent to the odd-dimensionality of the associated $H$-fixed point space $\mathcal{V}_{m,j}^{H}$, the full power of Theorem 7.1 can be demonstrated using a complete list of maximal elements in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some fixed $m>0$ together with computations of the relevant basic degrees $\deg_{\mathcal{V}_{m,j}}$ for $j\in\{0,1,2,3,4\}$. This can be achieved for $m=1$ (and then extended to any $m\in{\mathbb{N}}$ via $s$-folding, cf. Section 3.1) by the following GAP code:
1# A G.A.P. Program for the computation of Maximal Orbit Types and Basic Degrees associated with the G-isotypic decomposition V = V_0 $\times$ V_1 $\times$ V_2 $\times$ V_3 $\times$ V_4
2LoadPackage ("EquiDeg");
3# create groups O(2)xD8xZ2
4o2:=OrthogonalGroupOverReal(2);
5d8:=pDihedralGroup(8);
6z2:=pCyclicGroup(2);
7# generate D8xZ2
8g1:=DirectProduct(d8,z2);
9# set names for subgroup conjugacy classes in D8xZ2
10SetCCSsAbbrv(g1,["Z1","Z2","D1t","D1z","D1","Z1m","Z1p","D1zt","D1pt","D1p","D2","Z4","D2t","D2zt","D2d","Z4d","D2dt","D2z","Z2p","Z4p","D4dt","D2p","D4","D2pt","D4z","D4zt","Z8","Z8d","D4t","D4d","D4p","Z8p","D8","D4pt","D8d","D8z","D8dt","D8p"]);
11# generate O(2)xD8xZ2
12G:=DirectProduct(o2,g1);
13ccs:=ConjugacyClassesSubgroups(G);
14# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component
15irrs := Irr(G);
16# The D_8-isotypic components 0,1,2,3,4 are indexed in GAP as 1,9,11,13,7
17for i in [1,9,11,13,7] do
[…]
19 max_orbtyps := MaximalOrbitTypes( irrs[1,i] );
20 basic_deg := BasicDegree( irrs[1,i] );
21 View(basic_deg);
[…]
23 View(max_orbtyps);
24od;
[…]
In the GAP package EquiDeg, developed for equivariant degree computations in the Burnside Ring (cf. [26]), the conjugacy class of an amalgamated subgroup (cf. Section 4)
$H=K_{O}{}^{\varphi_{O}}\times^{\varphi_{\Gamma}}_{L}K_{\Gamma},$
is identified by the quadruple $(K_{1},K_{2},Z_{1},Z_{2})$ where $Z_{1}:=\text{\rm Ker\,}\varphi_{1}$ and $Z_{2}:=\text{\rm Ker\,}\varphi_{2}$ with the notation
$(H)=:(K_{1}{}^{Z_{1}}\times^{Z_{2}}K_{2}).$
Using this modified amalgamated notation, we can express the output of the above program, i.e. the maximal elements in the isotropy lattices $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for $j\in\{0,1,2,3,4\}$ and the corresponding basic degrees as follows:
$\mathfrak{M}_{m,0}$ =$=\left\{(D_{m}\times D_{8}^{p})\right\};$
$\mathfrak{M}_{m,1}$ =$=\left\{(D_{4m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{8}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{d}}\tilde{D}_{4}^{p}),\;(D_{2m}{}^{D_{m}}\times^{D_{4}^{d}}D_{4}^{p})\right\};$
$\mathfrak{M}_{m,2}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{D_{2}^{d}}D_{2}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{2}^{d}}\tilde{D}_{2}^{p}),\;(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}D_{8}^{p})\right\};$
$\mathfrak{M}_{m,3}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{D_{1}^{p}}D_{2}^{p}),\;(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{1}^{p}}\tilde{D}_{2}^{p}),\;(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{8}^{p})\right\};$
$\mathfrak{M}_{m,4}$ =$=\left\{(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{p}}D_{8}^{p})\right\},$
and
$\deg_{\mathcal{V}_{m,0}}$ $=(G)-(D_{m}\times D_{8}^{p});$
$\deg_{\mathcal{V}_{1,1}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}\tilde{D}_{4}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{4}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{4}^{d}}{\mathbb{Z}}_{4}^{p})$
$\quad\quad-2(D_{4m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{4}^{d}}D_{8}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{d}}\tilde{D}_{4}^{p})-(D_{2m}{}^{D_{m}}\times^{D_{4}^{d}}D_{4}^{p});$
$\deg_{\mathcal{V}_{1,2}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}D_{2}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{2}^{-}}\tilde{D}_{2}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{2}^{-}}{\mathbb{Z}}_{2}^{p})$
$\quad\quad-(D_{2m}{}^{D_{m}}\times^{D_{2}^{d}}D_{2}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{2}^{d}}\tilde{D}_{2}^{p})-2(D_{8m}{}^{{\mathbb{Z}}_{m}}\times D_{8}^{p});$
$\deg_{\mathcal{V}_{1,3}}$ $=(G)+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{2}^{p})+2(D_{2m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}\tilde{D}_{2}^{p})+(D_{2m}{}^{D_{m}}\times^{{\mathbb{Z}}_{1}^{p}}{\mathbb{Z}}_{2}^{p})$
$\quad\quad-(D_{2m}{}^{D_{m}}\times^{D_{1}^{p}}D_{2}^{p})-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{1}^{d}}\tilde{D}_{2}^{p})-2(D_{8m}{}^{{\mathbb{Z}}_{m}}\times^{{\mathbb{Z}}_{1}^{p}}D_{8}^{p});$
$\deg_{\mathcal{V}_{m,0}}$ $=(G)-(D_{2m}{}^{D_{m}}\times^{\tilde{D}_{4}^{p}}D_{8}^{p}).$
In addition to the notation defined above, we have adopted the following ancillary shorthand for identification of subgroups in $D_{8}$ generated by the rotation $\gamma:=e^{\frac{2\pi i}{8}}$ and the relection $\kappa$:
${\mathbb{Z}}_{1}:=\{1\},\quad{\mathbb{Z}}_{2}:=\{1,\gamma^{4}\},\quad{\mathbb{Z}}_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6}\},$
$D_{1}:=\{1,\kappa\},\quad\tilde{D}_{1}:=\{1,\kappa\gamma\},\quad D_{2}:=\{1,\gamma^{4},\kappa,\kappa\gamma^{4}\},$
$\tilde{D}_{2}:=\{1,\gamma^{4},\kappa\gamma,\kappa\gamma^{5}\},\quad D_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6},\kappa,\kappa\gamma^{2},\kappa\gamma^{4},\kappa\gamma^{6}\},$
$\tilde{D}_{4}:=\{1,\gamma^{2},\gamma^{4},\gamma^{6},\kappa\gamma,\kappa\gamma^{3},\kappa\gamma^{5},\kappa\gamma^{7}\},$
$D_{8}:=\{1,\gamma,\gamma^{2},\gamma^{3},\gamma^{4},\gamma^{5},\gamma^{6},\gamma^{7},\kappa,\kappa\gamma,\kappa\gamma^{2},\kappa\gamma^{3},\kappa\gamma^{4},\kappa\gamma^{5},\kappa\gamma^{6},\kappa\gamma^{7}\}$
Likewise, the reader can consult the following list of shorthand to identify the subgroups in $D_{8}\times\mathbb{Z}_{2}$:
${\mathbb{Z}}_{1}^{p}=\{(1,1),(1,-1)\},\quad{\mathbb{Z}}_{2}^{-}=\{(1,1),(\gamma^{4},-1)\},\quad{\mathbb{Z}}_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1)\},$
$D_{1}^{p}:=\{(1,1),(\kappa,1),(1,-1),(\kappa,-1)\},\quad\tilde{D}_{1}^{p}:=\{(1,1),(\kappa\gamma,1),(1,-1),(\kappa\gamma,-1)\},$
$D_{2}^{d}:=\{(1,1),(\gamma^{4},-1),(\kappa,1),(\kappa\gamma^{4},-1)\},\quad\tilde{D}_{2}^{d}:=\{(1,1),(\gamma^{4},-1),(\kappa\gamma,1),(\kappa\gamma^{5},-1)\},$
$D_{2}^{p}:=\{(1,1),(\gamma^{4},1),(\kappa,1),(\kappa\gamma^{4},1),(1,-1),(\gamma^{4},-1),(\kappa,-1),(\kappa\gamma^{4},-1)\},$
$\tilde{D}_{2}^{p}:=\{(1,1),(\gamma^{4},1),(\kappa\gamma,1),(\kappa\gamma^{5},1),(1,-1),(\gamma^{4},-1),(\kappa\gamma,-1),(\kappa\gamma^{5},-1)\},$
$D_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1),(\kappa,1),(\kappa\gamma^{2},-1),(\kappa\gamma^{4},1),(\kappa\gamma^{6},-1)\}$
$\tilde{D}_{4}^{d}:=\{(1,1),(\gamma^{2},-1),(\gamma^{4},1),(\gamma^{6},-1),(\kappa\gamma,1),(\kappa\gamma^{3},-1),(\kappa\gamma^{5},1),(\kappa\gamma^{7},-1)\}$
$\tilde{D}_{4}^{p}:=\{(1,1),(\gamma^{2},1),(\gamma^{4},1),(\gamma^{6},1),(\kappa\gamma,1),(\kappa\gamma^{3},1),(\kappa\gamma^{5},1),(\kappa\gamma^{7},1),$
$\quad\quad\quad\quad(1,-1),(\gamma^{2},-1),(\gamma^{4},-1),(\gamma^{6},-1),(\kappa\gamma,-1),(\kappa\gamma^{3},-1),(\kappa\gamma^{5},-1),(\kappa\gamma^{7},-1),\}$
$D_{8}^{p}:=\{(1,1),(\gamma,1),(\gamma^{2},1),(\gamma^{3},1),(\gamma^{4},1),(\gamma^{5},1),(\gamma^{6},1),(\gamma^{7},1),(\kappa,1),(\kappa\gamma,1),(\kappa\gamma^{2},1),$
$\quad\quad\quad\quad(\kappa\gamma^{3},1),(\kappa\gamma^{4},1),(\kappa\gamma^{5},1),(\kappa\gamma^{6},1),(\kappa\gamma^{7},1),(1,-1),(\gamma,-1),(\gamma^{2},-1),(\gamma^{3},-1),$
$\quad\quad\quad\quad(\gamma^{6},-1),(\gamma^{7},-1),(\kappa,-1),(\kappa\gamma,-1),(\kappa\gamma^{2},-1),(\kappa\gamma^{3},-1),(\kappa\gamma^{4},-1),(\kappa\gamma^{5},-1),$
$\quad\quad\quad\quad(\kappa\gamma^{6},-1),(\gamma^{4},-1),(\gamma^{5},-1),(\kappa\gamma^{7},-1)\}$
At this point, Proposition 1.4 follows directly from the observations $(\rm i)$ that the fixed point set $\mathcal{V}_{m,j}^{H}$ is odd dimensional for any $(H)\in\mathfrak{M}_{m,j}$ and $(\rm ii)$ that the sets $\mathfrak{M}_{m,j}$, $m>0$, $j\in\{0,1,2,3,4\}$ are pairwise disjoint.